# Model Diffusion for Certifiable Few-shot Transfer Learning

### v1

@@ Abstract
In modern large-scale deep learning, a prevalent and effective workflow for solving low-data problems is adapting powerful pre-trained foundation models (FMs) to new tasks via parameter-efficient fine-tuning (PEFT). However, while empirically effective, the resulting solutions lack generalisation guarantees to certify their accuracy - which may be required for ethical or legal reasons prior to deployment in high-importance applications. In this paper we develop a novel transfer learning approach that is designed to facilitate non-vacuous learning theoretic generalisation guarantees for downstream tasks, even in the low-shot regime. Specifically, we first use upstream tasks to train a distribution over PEFT parameters. We then learn the downstream task by a sample-and-evaluate procedure – sampling plausible PEFTs from the trained diffusion model and selecting the one with the highest likelihood on the downstream data. Crucially, this confines our model hypothesis to a finite set of PEFT samples. In contrast to learning in the typical continuous hypothesis spaces of neural network weights, this facilitates tighter risk certificates. We instantiate our bound and show non-trivial generalization guarantees compared to existing learning approaches which lead to vacuous bounds in the low-shot regime.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Generalisation bounds for adapting CLIP to novel tasks (5-way classification with from 1 to 16 examples per class). Plots show error (y-axis) vs log-complexity term of the bound of Equation 3 (x-axis). Top/Bottom: The mean support/query (train/test) error for new tasks. The unshaded area corresponds to non-vacuous guarantees, while in the shaded area the total sum of both axes is vacuous. Other approaches are not even close to providing non-vacuous guarantees, while our approach can do so without sacrificing too much fit quality (top), or empirical test accuracy (bottom).
Generalisation certificates are crucial for high-importance applications where accuracy should be guaranteed for legal or ethical reasons. Guarantees should certify the minimum testing accuracy expected on unseen data drawn from the training distribution. However, it is hard to establish non-trivial guarantees for large neural networks, since large learning capacity tends to produce looser guarantees. As such, there have only been a few successful demonstrations of non-vacuous guarantees for contemporary neural networks, even in the large-data regime (Dziugaite & Roy, 2017; Perez-Ortiz et al., 2021; Lotfi et al., 2024).
What about learning with sparse rather than large data? The problem of low-data learning is highly topical, due to the plethora of important limited-data applications (Wang et al., 2020), but challenging due to the difficulty of learning a large number neural network parameters without overfitting. This need has inspired several lines of research that make use of different forms of knowledge transfer, including meta-learning (Hospedales et al., 2021) and parameter-efficient transfer learning (PEFT) (Hu et al., 2021) from foundation models. While PEFT methods have recently been more empirically effective, neither family of approach has produced methods that can provide low-shot generalisation guarantees, to our knowledge. From a learning theoretic perspective this is because existing algorithms still search a hypothesis space (e.g., all neural network weights $\theta\in\mathcal{R}^{N}$) large enough to make known bounds vacuous when instantiated.
@@
More specifically, in the upstream phase, we fit PEFT modules to available source tasks, and then train a parameter diffusion model to generate PEFTs according to this task distribution. In the downstream phase, we learn by sample-then-evaluate instead of traditional gradient descent. PEFT modules, unconditionally generated by the diffusion model, are scored using the target task training set, and learning is to choose the highest scoring module. Compared to the original set of upstream models, the diffusion model can be more compact, and can interpolate between the original model set to achieve higher accuracy. This procedure is gradient-free, which has some scalability benefits (Malladi et al., 2023; Rezk et al., 2024), but more importantly it facilitates the use of PAC-Bayes finite-hypothesis bounds to provide non-vacuous guarantees, all while maintaining similar empirical accuracy to mainstream few-shot learning approaches. Figure 1 shows some illustrative results, demonstrating our learner’s ability to maintain practical efficacy while being constrained to low enough complexity to provide non-vacuous guarantees (white zone).
@@ 2 Risk Certificates for Deep Models
Certifying the generalization performance of a prediction model has been a central topic and studied rigorously in theoretical machine learning (Vapnik, 1995; Shalev-Shwartz & Ben-David, 2014; Mohri et al., 2018). The popular Vapnik–Chervonenkis (VC), Rademacher, and PAC-Bayes bounds all aim to relate the empirical risk (which is computable) to the generalization risk (which is impossible to compute) via some inequality relations. In this section we discuss, with less mathematical rigor, the key ideas of risk certificates and main bottlenecks that hinder naive application of the existing learning theory methods to deep models.
Formally, let $h\in\mathcal{H}$ be a hypothesis, a prediction function $y=h(x)$ that returns output $y$ (e.g., class label) for a given input $x$, and let $\mathcal{H}$ be the hypothesis space, a set of hypotheses from which a learning algorithm can select a hypothesis. In modern deep learning practice $h$ corresponds to a deep model with parameters (weights and biases) $\theta$, where we denote by $\Theta$ ($\ni\theta$) the set of all possible parameter values that the deep model can take. So $\Theta$ can be seen as the hypothesis space. A learning algorithm (e.g., SGD) is a specific recipe to choose $\theta$ from $\Theta$ for a given empirical data $S=\{(x_{i},y_{i})\}_{i=1}^{n}$ which is assumed to be i.i.d. samples from some underlying but unknown distribution $T$.
The ultimate goal is to find a hypothesis $\theta$ that minimizes the generalization error or risk $R(\theta)=\mathbb{E}_{(x,y)\sim T}[l(\theta;x,y)]$ where $l(\theta;x,y)$ is the instance risk (error or loss) (e.g., the cross-entropy loss or 0/1 error) of the hypothesis $\theta$ on the instance $(x,y)$. Since it is not possible to access the population $T$ in the expectation, even computing $R(\theta)$ is impossible. Hence one can rather minimize the empirical risk $r(\theta)=\frac{1}{n}\sum_{(x,y)\in S}l(\theta;x,y)$, instead. But $r(\theta)$ is just a surrogate for $R(\theta)$, and it does not give any certificate about the true generalization risk. In theoretical machine learning, several theorems have been proposed to relate $R$ and $r$, mostly via upper bounding $R$ in terms of $r$. That is, they typically follow the following form. For any $\theta\in\Theta$:
@@
where the complexity term is determined by the empirical data size $n=|S|$ and the so-called the hypothesis space complexity $\dim(\Theta)$ which quantifies how large/small your hypothesis space is. Roughly saying, the complexity term decreases as $n$ goes large and $\dim({\Theta})$ becomes small. The well-known VC, Rademacher, and PAC-Bayes bounds all follow the above form111 In the PAC-Bayes bounds, the risks are measured as expected risks over some (posterior) distribution over $\theta$ rather than point estimates. Also the complexity term is expressed in terms of divergence from a prior distribution. But if we confine the posterior to be sharply concentrated at a single point $\theta$ and use a flat prior, this roughly follows the form of (1). Also certain PAC-Bayes bounds have nonlinear relation between $R$ and $r$, however, they can be approximated as (1), where this simplification does not affect the reasoning in our paper. , while they are differentiated (roughly) by how the complexity term is defined precisely.
Note that the right hand side of (1) is always computable (both $r$ and the complexity term), and called the risk certificate. It is called so since the generalization risk is guaranteed to be no greater than the computed value (i.e., upper bounding). However, when the computed risk certificate is greater than 1, assuming bounded loss $l\in[0,1]$, then it is said vacuous since saying the generalization risk is less than 1 is useless. However, if it is less than 1 (or sometimes less than the random guess risk in the classification setting), then we say that it is non-vacuous. Several traditional models and algorithms (e.g., linear SVM) were shown to have non-vacuous risk certificates when proper regularization is imposed to control the complexity term (Vapnik, 1995; Shalev-Shwartz & Ben-David, 2014).
Then the follow-up question is whether or not deep neural networks conventionally trained by gradient descent have non-vacuous bounds. The answer is no if we apply the traditional learning theory techniques directly, unless the data size $n$ is huge. It is mainly due to the large number of parameters in a deep model that leads to high $\dim(\Theta)$. Although recent sparse adapter (aka PEFT) tricks in deep learning (e.g., LoRA (Hu et al., 2021) and LoRA-XS (Balazy et al., 2024)) can be considered, the gradient-based learning still suffers from the large hypothesis space complexity due to the inherently continuous space of $\Theta$. To our knowledge, no existing deep learning approach can achieve meaningful generalization bounds in the low-shot regime.
@@ 3.1 Our Approach at High Level
In this paper, we approach the problem in a completely different perspective. Our key idea is to avoid the continuous hypothesis space and gradient-based learning algorithms by devising a finite hypothesis space and adopting an (gradient-free) evaluate-then-select learning algorithm. First, we focus on the multi-task transfer learning setup where a large number of training tasks are available. We learn a task distribution from these tasks using a diffusion generative model, where we build a finite hypothesis space $\Theta$ comprised of a large number of samples $\theta$ generated from the learned diffusion model. The diffusion models are known to have strong interpolation capability in density estimation, and our hope is that the constructed $\Theta$ is large enough (but still finite) to cover any downstream tasks successfully as long as they come from the same underlying task distribution.
Our choice of learning algorithm is as simple as: evaluate each $\theta\in\Theta$ and choose the one with the minimum empirical risk. In the large cardinality $\Theta$ cases, we make use of efficient and smart heuristic search algorithms to reduce the forward-pass overhead (as detailed in Sec. 3). Our diffusion-based finite hypothesis space and the gradient-free learning algorithm, when combined with existing finite-hypothesis PAC-Bayes bounds, surprisingly leads to tight non-vacuous risk certificates on large-scale LLM/vision benchmarks with large-scale FLAN-T5/CLIP models. In addition the actual test performance of our approach on these benchmarks is comparable to standard learning algorithms. We do not pay a huge price for these guarantees. We also note that our evaluate-then-select learning algorithm essentially performs PAC-Bayes bound minimization since the complexity term is fixed/constant (details in Sec. 3). This turns out to be the optimal learning strategy in terms of risk certificates.
Before we jump into our main approach in greater detail in Sec. 3.3, we will formally define the problem setup.
@@ 3.2 Problem Setup and Notation
We describe the low-shot cross-task transfer learning problem that we aim to solve in the paper. Consider that we have a training pool of related tasks $T_{1},\dots,T_{N}$, i.i.d. sampled from some unknown but true task distribution $p_{true}(T)$. At test time, we are given a new test task $T^{*}$ sampled from the same $p_{true}(T)$, but only in the form of low-shot data samples $S^{*}=\{(x_{i},x_{y})\}_{i=1}^{n}$ (aka support data) from $T^{*}$. By low shot, we mean the number of support samples $n$ is small. Our goal is to find a way to ensure tight generalization bounds for the underlying deep models. As discussed in Sec. 2, a main challenge here is that we have low-shot data $S^{*}$ and a large number of model parameters, where the latter immediately translates into high hypothesis space complexity. Hence, applying the traditional learning theories directly to this problem leads to vacuous error bounds. We come up with a new method that exploits the training tasks to transfer the knowledge to unseen tasks so that it can offer tight non-vacuous risk certificates.
3.3 Details of Our Approach
@@
Our first observation is that the gradient-based model adaptation to low-shot data samples $S^{*}$ needs to avoided to reduce the hypothesis space complexity as discussed in Sec. 2. Our key intuition is that we can learn the task distribution $p_{true}(T)$ using the training tasks $\{T_{i}\}_{i=1}^{N}$, but while doing so, we will introduce some strong inductive bias or regularization. Let $\theta_{i}$ be the learned neural network parameters for task $T_{i}$. (Throughout the paper, we deal with PEFT adapter parameters as $\theta$, leaving the pre-trained main backbone parameters fixed.) We can view $\theta_{i}$ as the best description for the task $T_{i}$. We can collect task-wise trained network parameters $\{\theta_{i}\}_{i=1}^{N}$, and learning the task distribution $p_{true}(T)$ can be regarded as a density estimation problem, namely estimating the density $p(\theta)$ from the i.i.d. samples $\{\theta_{i}\}_{i=1}^{N}$. In other words, we treat $p(\theta)$ as a parametric description or surrogate for $p_{true}(T)$.
@@
We learn a diffusion model $p(\theta)$ with $\{\theta_{i}\}_{i=1}^{N}$ as training data following (Ho et al., 2020). The estimated $p(\theta)$, as a proxy for $p_{true}(T)$, can be used at test time: We generate many plausible candidate samples $\theta$ (i.e., tasks $T$), among which we select the one that is closest to $T^{*}$. Since we only have the support data set $S^{*}$ of $T^{*}$ available, we find the candidate sample that has the least discrepancy with $S^{*}$. The least discrepancy criterion can be defined as the minimum loss selection rule, i.e., $\arg\min_{\theta\in\Theta}\sum_{(x,y)\in S^{*}}l(\theta;x,y)$ where $\Theta$ is the set of candidate samples from the diffusion model $p(\theta)$. Importantly, this is the empirical risk minimization where the candidate diffusion sample set $\Theta$ serves as our hypothesis space.
@@
It is important to note that this learning strategy is nothing but selection from a finite hypothesis space, instead of searching a continuous hypothesis space as in gradient-based parameter-level fine-tuning. The choice of the finite hypothesis set (diffusion model samples) prior to learning provides an opportunity for a strong regularizing inductive bias that can be learned from the upstream tasks (by training the diffusion model).
@@ 3.4 Risk Certificate with Finite Hypothesis Space
Suppose we have a well-trained zoo of (PEFT adapter) parameters $\overline{\Theta}=\{\theta_{i}\}_{i=1}^{N}$, where each $\theta_{i}$ is the optimal parameters for the $i$-th training task ($i=1,\dots,N$). Without the diffusion training with $\overline{\Theta}$, nothing can prevent us from using the the model zoo $\overline{\Theta}$ itself as our hypothesis space, i.e., $\Theta=\overline{\Theta}$. This becomes more reasonable as $N$ goes larger since $\overline{\Theta}$ would be richer and closer to the true $p_{true}(T)$. We will call this strategy of $\Theta=\overline{\Theta}$ the model zoo strategy. In contrast, another proposal of ours is to train a diffusion model $p(\theta)$ with $\overline{\Theta}$, and build $\Theta$ using the samples from it, which we call the STEEL [^2] Alternatively, we can union model zoo and diffusion samples together to build a larger hypothesis space. This is practically effective, but in this paper we focus only on model zoo and diffusion strategies so as to contrast their behaviors more carefully.. Whereas both are our proposals, the diffusion is our main strategy as it is more attractive in the following aspects: i) the model zoo strategy requires a large amount of space to store the $N$ models while the diffusion strategy is more scalable, only needing to store the diffusion model itself; ii) The diffusion model is widely known to have strong interpolation capability to approximate the target density better, which in practice can provide improved test accuracy
@@
A crucial benefit of our test-time adaptation strategy (2) is that we have a tight provable generalization error bound that can serve as a risk certificate for its test-time prediction quality. This mainly originates from the finite hypothesis space $\Theta$. More specifically, using the PAC-Bayes theorems (e.g., Sec. 2.1.3 in (Alquier, 2021)), we can show that with probability at least $1-\epsilon$,
@@
However, the computational question naturally arises: How do we solve (2) efficiently? We consider two solutions:.
@@
We emphasize again that in all these three strategies, the generalization error bound (3) holds true, but with possibly differently/suboptimally selected $\theta$s in the latter case, which may imply (slight) increase in the empirical loss $r(\theta)$, and hence a slight loosening of the obtained certificate.

### v2

@@ Abstract
In contemporary deep learning, a prevalent and effective workflow for solving low-data problems is adapting powerful pre-trained foundation models (FMs) to new tasks via parameter-efficient fine-tuning (PEFT). However, while empirically effective, the resulting solutions lack generalisation guarantees to certify their accuracy - which may be required for ethical or legal reasons prior to deployment in high-importance applications. In this paper we develop a novel transfer learning approach that is designed to facilitate non-vacuous learning theoretic generalisation guarantees for downstream tasks, even in the low-shot regime. Specifically, we first use upstream tasks to train a distribution over PEFT parameters. We then learn the downstream task by a sample-and-evaluate procedure – sampling plausible PEFTs from the trained diffusion model and selecting the one with the highest likelihood on the downstream data. Crucially, this confines our model hypothesis to a finite set of PEFT samples. In contrast to the typical continuous hypothesis spaces of neural network weights, this facilitates tighter risk certificates. We instantiate our bound and show non-trivial generalization guarantees compared to existing learning approaches which lead to vacuous bounds in the low-shot regime.
@@ 1 Introduction
Figure 1: Generalization bounds for adapting CLIP to novel tasks (5-way classification with 1–16 examples per class). Plots show classification error (y-axis) versus the complexity term (x-axis, log scale; square root terms from Equations 3, and 7). Top/Bottom: Mean support/query (train/test) error on new tasks. Shaded regions indicate vacuous bounds, where (support error + complexity) $\geq 1$. Non-vacuous guarantees lie in the unshaded region. Competing methods (SGD, BBPT) fail to achieve non-vacuous bounds. In contrast, our method yields non-vacuous guarantees without significantly compromising training fit (top) or test accuracy (bottom).
Generalisation certificates are crucial for high-importance applications where accuracy should be guaranteed for legal or ethical reasons. Guarantees should certify the minimum testing accuracy expected on unseen data drawn from the training distribution. However, it is hard to establish non-trivial guarantees for large neural networks, since large learning capacity tends to produce looser guarantees. As such, there have only been a few successful demonstrations of non-vacuous guarantees for contemporary neural networks, even in the large-data regime [7, 21, 15].
What about learning with sparse rather than large data? The problem of low-data learning is highly topical, due to the plethora of important limited-data applications [36], but challenging due to the difficulty of learning a large number neural network parameters without overfitting. This need has inspired several lines of research that make use of different forms of knowledge transfer, including meta-learning [10] and parameter-efficient transfer learning (PEFT) [11] from foundation models. While PEFT methods have recently been more empirically effective, neither family of approach has produced methods that can provide low-shot generalisation guarantees, to our knowledge. From a learning theoretic perspective this is because existing algorithms still search a hypothesis space (e.g., all neural network weights $\theta\in\mathcal{R}^{N}$) large enough to make known bounds vacuous when instantiated.
@@
More specifically, in the upstream phase, we fit PEFT modules to available source tasks, and then train a parameter diffusion model to generate PEFTs according to this task distribution. In the downstream phase, we learn by sample-then-evaluate instead of traditional gradient descent. PEFT modules, unconditionally generated by the diffusion model, are scored using the target task training set, and learning is to choose the highest scoring module. Compared to the original set of upstream models, the diffusion model can be more compact, and can interpolate between the original model set to achieve higher accuracy. This procedure is gradient-free, which has some scalability benefits [17, 23], but more importantly it facilitates the use of PAC-Bayes finite-hypothesis bounds to provide non-vacuous guarantees, all while maintaining similar empirical accuracy to mainstream few-shot learning approaches. Figure 1 shows some illustrative results, demonstrating our learner’s ability to maintain practical efficacy while being constrained to low enough complexity to provide non-vacuous guarantees (white zone).
@@ 2 Risk Certificates for Deep Models
Certifying model generalization performance is fundamental in theoretical machine learning [32, 26, 19]. Vapnik–Chervonenkis (VC), Rademacher, and PAC-Bayes bounds connect empirical risk (computable) to generalization risk (impossible to compute) through inequalities. Here we discuss the core concepts of risk certificates and why they are challenging to apply to deep models.
Let $h\in\mathcal{H}$ be a hypothesis (prediction function $y=h(x)$) and $\mathcal{H}$ the hypothesis space. In deep learning, $h$ corresponds to a model with parameters $\theta$, and $\Theta$ ($\ni\theta$) represents all possible parameter values, serving as the hypothesis space. A learning algorithm (i.e, SGD) selects $\theta$ from $\Theta$ given empirical data $S={(x_{i},y_{i})}_{i=1}^{n}$ sampled i.i.d. from distribution $T$. The goal is minimizing generalization risk $R(\theta)=\mathbb{E}_{(x,y)\sim T}[l(\theta;x,y)]$, where $l(\theta;x,y)$ is the instance risk. Since computing $R(\theta)$ is impossible without access to $T$, we minimize empirical risk $r(\theta)=\frac{1}{n}\sum_{(x,y)\in S}l(\theta;x,y)$. Nevertheless, $r(\theta)$ is a surrogate for $R(\theta)$ and does not give any certificate about true generalization risk. Therefore, theoretical bounds relate $R$ and $r$ as follows. For any $\theta\in\Theta$:
@@
where the complexity term depends on data size $n$ and hypothesis space complexity $\dim(\Theta)$, decreasing as $n$ increases and $\dim({\Theta})$ decreases111 In the PAC-Bayes bounds, the risks are measured as expected risks over some (posterior) distribution over $\theta$ rather than point estimates. Also the complexity term is expressed in terms of divergence from a prior distribution. But if we confine the posterior to be sharply concentrated at a single point $\theta$ and use a flat prior, this roughly follows the form of (1). Also certain PAC-Bayes bounds have nonlinear relation between $R$ and $r$, however, they can be approximated as (1), where this simplification does not affect the reasoning in our paper. . The right side of (1) is the risk certificate, which guarantees an upper bound on the risk of generalization. All terms on the right side are also computable. Different bounding methods, for example VC and Pac-Bayes bounds, produce different complexity terms. In totality, certificates above 1 (with $l\in[0,1]$) are vacuous, while we refer to those below 1 (or sometimes less than the random guess risk in the classification setting) as non-vacuous.
Traditional models like linear SVM achieve non-vacuous bounds through proper regularization [32, 26]. However, conventional deep neural networks trained by gradient descent lack non-vacuous bounds unless $n$ is extremely large, due to high $\dim(\Theta)$ from numerous continuous parameters. Even sparse adapter methods (e.g., LoRA [11], LoRA-XS [3]) face challenges from the continuous nature of $\Theta$. No existing deep learning approach achieves meaningful generalization bounds in low-shot settings.
@@ 3.1 Our Approach at High Level
We propose a novel approach using a finite hypothesis space and gradient-free learning, departing from traditional continuous methods. In multi-task transfer learning, we learn a task distribution via a diffusion generative model, creating a finite hypothesis space $\Theta$ from model-generated samples. Our simple learning algorithm selects the $\theta\in\Theta$ with minimal empirical risk, using heuristic search for efficiency with large $\Theta$ to reduce the forward-pass overhead (as detailed in Sec. 3).
This approach, combined with finite-hypothesis PAC-Bayes bounds, yields tight non-vacuous risk certificates on large-scale LLM/vision benchmarks using FLAN-T5/CLIP models. Importantly, test performance remains comparable to standard learning algorithms. The evaluate-then-select strategy optimizes PAC-Bayes bounds by keeping the complexity term constant. We present the formal problem setup below before detailing our approach in Sec. 3.3.
@@ 3.2 Problem Setup and Notation
We describe the low-shot cross-task transfer learning problem that we aim to solve in the paper. Consider that we have a training pool of related tasks $T_{1},\dots,T_{N}$, i.i.d. sampled from some unknown but true task distribution $p_{true}(T)$. At test time, we are given a new test task $T^{*}$ sampled from the same $p_{true}(T)$, but only in the form of low-shot data samples $S^{*}=\{(x_{i},x_{y})\}_{i=1}^{n}$ (aka support data) from $T^{*}$. By low shot, we mean the number of support samples $n$ is small. We remark that in terms of problem specification and assumptions, this is almost exactly the same as assumed by both the meta-learning [10] and model-zoo [12] literature.
Our goal is to find a way to ensure tight generalization bounds for the underlying deep models at test time. As discussed in Sec. 2, a main challenge here is that we have low-shot data $S^{*}$ and a large number of model parameters, where the latter immediately translates into high hypothesis space complexity. Hence, applying the traditional learning theories directly to this problem leads to vacuous error bounds. We come up with a new method that exploits the training tasks to transfer the knowledge to unseen tasks so that it can offer tight non-vacuous risk certificates.
3.3 Transfer Learning by Sampling and Evaluation
@@
Our first observation is that gradient-based model adaptation to low-shot samples $S^{*}$ must be avoided to reduce hypothesis space complexity (Sec. 2). Our key intuition is that we can learn the task distribution $p_{true}(T)$ from training tasks $\{T_{i}\}_{i=1}^{N}$, but doing so introduces a strong inductive bias or regularization. Let $\theta_{i}$ be the learned neural network parameters for task $T_{i}$. (Throughout, we treat PEFT adapter parameters as $\theta$, keeping the pre-trained backbone fixed.) We view $\theta_{i}$ as the best description of $T_{i}$ and collect task-wise parameters $\{\theta_{i}\}_{i=1}^{N}$. Learning $p_{true}(T)$ thus reduces to a density estimation problem, i.e., estimating $p(\theta)$ from i.i.d. samples $\{\theta_{i}\}_{i=1}^{N}$, treating $p(\theta)$ as a surrogate for $p_{true}(T)$.
@@
We learn a diffusion model $p(\theta)$ with $\{\theta_{i}\}_{i=1}^{N}$ as training data following Ho et al. [8]. At test time, the estimated $p(\theta)$ serves as a proxy for $p_{true}(T)$. We generate plausible candidate samples $\theta$ (i.e., tasks $T$) and select the one closest to $T^{*}$. Since only $S^{*}$ of $T^{*}$ is available, we choose the sample with the least discrepancy from $S^{*}$ via the minimum loss rule: $\arg\min_{\theta\in\Theta}\sum_{(x,y)\in S^{*}}l(\theta;x,y)$, where $\Theta$ is the set of diffusion model candidates. This corresponds to empirical risk minimization where $\Theta$ acts as the hypothesis space, and we this learner STEEL (Sample ThEn Evaluate Learner).
@@
This strategy amounts to selection from a finite hypothesis space, rather than searching a continuous space via gradient-based fine-tuning. The choice of a finite hypothesis set (diffusion model samples) before learning enables a strong regularizing inductive bias learned from upstream tasks.
@@ 3.4 Risk Certificate with Finite Hypothesis Space
Suppose we have a well-trained zoo of (PEFT adapter) parameters $\overline{\Theta}=\{\theta_{i}\}_{i=1}^{N}$, where each $\theta_{i}$ is the optimal parameters for the $i$-th training task ($i=1,\dots,N$). Without the diffusion training with $\overline{\Theta}$, nothing can prevent us from using the the model zoo $\overline{\Theta}$ itself as our hypothesis space, i.e., $\Theta=\overline{\Theta}$. This becomes more reasonable as $N$ goes larger since $\overline{\Theta}$ would be richer and closer to the true $p_{true}(T)$. We will call this strategy of $\Theta=\overline{\Theta}$ the model zoo strategy. To go beyond this, our full STEEL method trains a diffusion model $p(\theta)$ with $\overline{\Theta}$, and builds $\Theta$ using the samples from it [^2] Alternatively, we can union model zoo and diffusion samples together to build a larger hypothesis space. This is practically effective, but in this paper we focus only on model zoo and diffusion strategies so as to contrast their behaviors more carefully.. Whereas both are our proposals, the diffusion is our main strategy as it is more attractive in the following aspects: i) the model zoo strategy requires a large amount of space to store the $N$ models while the diffusion strategy is more scalable, only needing to store the diffusion model itself; ii) The diffusion model is widely known to have strong interpolation capability to approximate the target density better, which in practice can provide improved test accuracy
@@
A crucial benefit of our test-time adaptation strategy (2) is that we have a tight provable generalization error bound that can serve as a risk certificate for its test-time prediction quality. This mainly originates from the finite hypothesis space $\Theta$. More specifically, using the PAC-Bayes theorems (e.g., Sec. 2.1.3 in [1]), we can show that with probability at least $1-\epsilon$,
@@
However, the computational question naturally arises: How do we solve (2) efficiently? We consider two solutions:
@@
We emphasize again that in all these strategies, the generalization error bound (3) holds true, but with possibly differently/suboptimally selected $\theta$s in the latter case, which may imply (slight) increase in the empirical loss $r(\theta)$, and hence a slight loosening of the obtained certificate.